# Synaptic Pruning: A Biological Inspiration for Deep Learning Regularization

### v1

@@ 6 Conclusion and Future Work
Future work should address the computational overhead through optimized implementation and parallelization techniques with validation on larger datasets. Long-term research directions should explore implementations that include both synaptic growth and pruning phases, potentially leading to next-generation bio-inspired architectures that fundamentally change how neural networks learn and adapt. By continuing to incorporate neuro-biological principles into deep learning, we envision improved and novel techniques that enable neural networks to better emulate the remarkable efficiency and generalization capabilities of biological learning systems.

### v2

@@ 6 Conclusion and Future Work
Future work should address the computational overhead through optimized implementation and parallelization techniques with validation on larger datasets. Long-term research directions should explore implementations that include both synaptic growth and pruning phases, potentially leading to next-generation bio-inspired architectures that fundamentally change how neural networks learn and adapt. By continuing to incorporate neuro-biological principles into deep learning, we envision improved and novel techniques that enable neural networks to better emulate the remarkable efficiency and generalization capabilities of biological learning systems. To support further research in this area, the full source code used in this study is available on GitHub at https://github.com/xalentis/SynapticPruning